# A Portable and Stealthy Inaudible Voice Attack Based on Acoustic Metamaterials

### v1

@@ IX RELATED WORK
IX-A Attacks on Voice Control Systems
Attacks on voice control systems have been proven to cause serious security issues. The previous attacks are mainly into two types: attacks based on adversarial examples and inaudible attacks. The attacks based on adversarial examples are presented by deceiving the victims [59, 60, 61]. However, these attacks based on adversarial examples rely on audible sounds, which can cause suspicion among alert victims. Inaudible attacks can directly disrupt or control voice control systems through ultrasound, which makes them more covert [7, 8, 9, 14]. For example, DolphinAttack [8] modulates attack commands onto ultrasound waves for attack, achieving control over smart devices. LipRead [9] and DolphinAttack (Long-range System) [15] expands the range of inaudible attacks on this basis, but its large devices are easily detected. Our MetaAttack first combines attacks with acoustic metamaterials to achieve long-range attacks in a compact size, making it more covert compared to previous studies.
IX-B Acoustic Metamaterials
Acoustic metamaterial designed for ventilation barriers is an emerging research field. This acoustic metamaterial can achieve the filtering function while maintaining ventilation through compact size. For example, Ghaffarivardavagh et al. [42] proposed a deep subwavelength acoustic metamaterial using Fano-like resonance. Subsequently, Zhu et al. [25] increased the filtering bandwidth of the acoustic metamaterial through nonlocal effects (600 Hz - 1900 Hz). Based on the above research, MetaAttack further expands the filtering range of this type of metamaterials (100 Hz - 4000 Hz) and first expands its application field to the inaudible attacks.

### v2

@@ IX RELATED WORK
Attacks on Voice Control Systems. Attacks on voice control systems have been proven to pose serious security threats. Existing attack methods mainly fall into the following categories: adversarial attacks, magnetic injection attacks, laser attacks, and inaudible attacks.
Adversarial attacks deceive devices by embedding carefully crafted instructions in normal audio to induce them to perform actions preset by the attacker [67, 68, 69, 70, 71, 72, 73, 74, 75]. For example, CommanderSong [67] hides commands within songs and performs the attack by playing the song. Metamorph [68] enhances robustness using pre-coding, extending the attack range to 6 meters. In addition, MagBackdoor [17] injects command signals into voice control systems through the device’s charging port, while Sugawara et al. [76] use long-range lasers to simulate sound vibrations to attack voice control systems. Inaudible attacks can silently interfere with or control voice control systems, making them more covert [11, 12, 13, 19, 42, 20, 18, 46]. For example, DolphinAttack [12] achieves control over smart devices by modulating attack commands onto ultrasound waves. LipRead [13] and DolphinAttack (Long-range System) [20] further extend the attack range by incorporating spectrum splitting techniques and large speaker arrays. VRIFLE [18] injects ultrasonic signals while the target is speaking, interfering with the target’s control of the device, thereby increasing the destructiveness of the attack. MetaAttack first combines attacks with acoustic metamaterials to achieve long-range attacks in a compact size, offering both concealment and threat.
Acoustic Metamaterials. Acoustic metamaterials designed for ventilation barriers are an emerging research area, enabling both filtering and airflow in a compact form. For example, Ghaffarivardavagh et al. [51] proposed a deep subwavelength acoustic metamaterial based on Fano-like resonance, and Zhu et al. [30] expanded its filtering bandwidth using nonlocal effects. Building on this, MetaAttack further extends the filtering range to 100 – 4000‘Hz and pioneers its application in ultrasonic attacks.
@@ X CONCLUSION
XI Ethical Considerations
Ethical Challenges in Reproducibility and Disclosure. Although publishing the 3D model designs of MetaAttack would aid in the reproducibility of the experiment, it could also provide malicious actors with the information necessary to execute attacks. Therefore, we have decided to share the related materials only with verified researchers who contact us, balancing transparency with security.
Ethical Considerations in Experiments and Testing. The research equipment was purchased at our own expense, and all experiments were approved by the university’s Institutional Review Board (IRB), ensuring compliance with data security and ethical standards. All participants voluntarily agreed to participate in the experiments based on informed consent, ensuring adherence to ethical requirements.